# New Approach to Strong Correlation: Twisting Hubbard into the Orbital Hatsugai-Kohmoto Model

### v2

@@ III Orbital HK
Figure 4: LESW and DSWT in the exact solution of the 4-orbital HK model ($U=10$) for the hopping parameters shown in panels (a) and (b) respectively. The solid line shown with slope $2x$ is the band HK or atomic Hubbard result. The dashed and dotted lines depict the semiconductor and Fermi liquid results respectively. Note, there is no qualitative difference with the exact diagonalization results for the Hubbard modelMeinders et al. (1993); Eskes et al. (1991). As $t/U$ increases, so does the DSWT.
All of these results point to a fundamental connection with the Hubbard model. Fig. 2 contains the essential idea. Because of the constraint, $N\times n=N_{\rm sites}$ and $\text{rBZ}=\text{BZ}/n$, $N$ the number of k-points, the rBZ is half the initial BZ when $n=2$, quartered for $n=4$, etc. as illustrated in Fig. 2a. Consequently, when $n=N_{\rm sites}$, no momentum degrees of freedom exist and only local orbitals are left, thereby leading to the $N_{\rm sites}$ Hubbard model. Alternatively, we can think of the orbital index as a label for the number of atoms per unit cell as depicted in (Fig. 2(b-d)). $N_{\rm sites}/n$, or $N$, is now the number of cells. Each cell contains the exact information of the $n$-site Hubbard model. For example, $n=2$ signifies that the orbital basis contains the information of the Hubbard dimer which has a non-degenerate ground state, and hence the spurious $2^{N}$ spin degeneracy of the band-HK model is removed. In addition to hybridization between orbitals in each cell, the cells are also connected in terms of hopping and interactions. For $n=4$, we show the explicit connection between the cells. The matrix $g({\bf k})$, representing the hopping connectivities in Fig. 2c, is given by
$\left(\begin{array}[]{cccc}0&\varepsilon_{tx}&\varepsilon_{ty}&\varepsilon_{t^{\prime}}\\ \varepsilon_{tx}&0&\varepsilon_{t^{\prime}}&\varepsilon_{ty}\\ \varepsilon_{ty}&\varepsilon_{t^{\prime}}&0&\varepsilon_{tx}\\ \varepsilon_{t^{\prime}}&\varepsilon_{ty}&\varepsilon_{tx}&0\end{array}\right)$ (9)
@@
for the 4-atom case, with $\varepsilon_{tx}=-2t\cos k_{x}$, $\varepsilon_{ty}=-2t\cos k_{y}$ and $\varepsilon_{t^{\prime}}=-4t^{\prime}\cos k_{x}\cos k_{y}$, identical in form to that in the $2\times 2$ Hubbard cluster with twisted boundary conditions (TBC) Huang (2020) where $\alpha$ is replaced with the site index $i$ and $k_{x}=\theta_{x}$ and $k_{y}=\theta_{y}$ where $\theta_{x}$ and $\theta_{y}$ define the TBC. Besides, TBC gives local-in-momentum interactions in the reduced BZ which is exactly the interaction term in Eq. (1). For any $n<N_{\rm sites}$, the orbital HK model corresponds to a $N_{\rm sites}/n$ copies of $n$-site Hubbard clusters all connected by TBC. The matrix that determines the TBC is purely local as it is determined by the dynamics. Hence, orbital HK is purely local in the rBZ regardless of the number of orbitals used, even $n=1$ corresponding to band HK whose locality is illustrated in Fig. 1. In this case, the TBC matrix is simply $\varepsilon_{\bf k}$. This exact mapping is a surprise and offers a new way of interpreting exact diagonalization with TBC. The most general statement is as follows. Theorem: The orbital HK model is equivalent to covering the Brillouin zone with $N_{\rm sites}/n$ Hubbard clusters each containing $n$-sites all connected via TBC. Clearly, the $N_{\rm sites}$-cluster Hubbard model results when $n=N_{\rm sites}$. Loosely, the $N_{\rm sites}/n$ identical copies throughout the Brillouin zone can be thought of as replica copies.
@@
Aside from DSWT, the orbital HK model also exhibits a pseudogap in the density of states. Shown in Fig. 5 is the density of states for the 4-orbital HK model with (Fig. 5(a,c)) and without (Fig. 5(b,d)) the next-nearest-neighbor hopping, $t^{\prime}$. Fig. 5(a) shows a suppression of the density of states at the chemical potential ($\rho(\omega=0)\lesssim 0.1$) in the underdoped region with $t^{\prime}=-0.25$ when $U\geq W$ ($W=8t$ is the bare bandwidth), indicative of a pseudogap in the absence of superconducting order. The complete tracking of the density of states at zero frequency in Fig. 5(c) displays the trend that as $U$ increases, the pseudo-gap region appears and extends to higher hole-doped density. In contrast, when $t^{\prime}=0$, the pseudo-gap suppression does not arise until $U\gg W$, as shown in Fig. 5(b,d). These observations are consistent with a recent studyHuang (2020) on a $2\times 2$ Hubbard cluster with TBC.
@@
Finally, we compute the heat capacity of the 4-orbital HK model at half-filling shown in Fig. 6. Most noticeable is the two-peak structure at $U>W$ representing a demarcation of the charge and spin degrees of freedom into high and low-temperature regimes, respectively, consistent with QMC simulations on the Hubbard model Duffy and Moreo (1997); Wang et al. (2022). Also at half-filling, we find a near-crossing of the heat capacity curves as a function of $U$ at a temperature intermediate between the spin and charge excitations. The Maxwell relations governing the entropy dictateVollhardt (1997) such a crossing. Since there is no sign-problem restriction, we can access the low-temperature heat capacity exactly. As we show in Fig. 6(b), the low-temperature heat capacity data follow a power-law increase detailed in the supplement. While such algebraic growth might seem counterintuitive for a Mott insulator, the first-excited state is charge neutralMai et al. (2024); Zhao et al. (2023a) as shown in the supplement. Such charge-neutral excitations determine the low-T behavior of the specific heat.
@@
IV Concluding Remarks
We have shown here that orbital HK offers a powerful analytical and numerical tool for treating Mott physics described by the Hubbard model. While the agreement with Hubbard is expected in the large orbital number limit, already at $n=2,4$, we find spot-on agreement with Hubbard for double occupancy at half-filling and $1/8$-hole-doping as Fig. (3c) attests. This is not a surprise. After all, we have shown that HK represents a fixed pointZhao et al. (2023b); Huang et al. (2022) for Mott physics in which not even Hubbard is a relevant perturbation. The fixed point obtains because HK contains the most relevant interaction with scaling dimension $-2$ of a Fermi liquid. The large negative scaling dimension stems from the single momentum integration. Adding orbitals to each ${\bf k}$-point does not lead to further momentum integrations as such decorations of each ${\bf k}$-point just appear as discrete summations. What the orbital content does add is non-commutativity between the kinetic and potential energy terms, thereby pushing it closer to the Hubbard model. Five manifestations of this non-commutativity stand out: 1) absence of a quasiparticle peak in the density of states at the Mott transition in agreement with standard Hubbard resultsMaier et al. (2005); Qin et al. (2016); LeBlanc et al. (2015), 2) the appearance of DSWT, 3) near-perfect quantitative agreement (see Fig. (3c)) with the QMC calculations on the Hubbard model for double occupancy that asymptotes to zero as $U$ increases, 4) a pseudogap in the density of states, and 5) separation of spin and charge excitations in the heat capacity into low and high-temperature regimes. As all of this physics is present in the Hubbard model, we conclude the orbital HK model is an effective simulator of Mott physics exemplified by the Hubbard model. This connection is apparent because the orbital extension is equivalent to replicas of finite Hubbard clusters covering the Brillouin zone all connected via TBC. Given the simplicity and cost-effective nature of our approach, orbital HK represents a new computational tool for strongly correlated Mott physics.
@@ IV Concluding Remarks
Acknowledgements We thank Shiwei Zhang for sending us his double occupancy data that is plotted in Fig. 3d (green dots) and George Sawatzky for a helpful e-mail exchange. This work was supported by the Center for Quantum Sensing and Quantum Materials, a DOE Energy Frontier Research Center, grant DE-SC0021238 (P. M. and P. W. P.). PWP also acknowledges NSF DMR-2111379 for partial funding of the HK work which led to these results.

### v3

@@ III Orbital HK
Figure 5: (a) double occupancy at half-filling ($x=0$) as a function of $U$ for the orbital HK model with various $n$. The green and gray dots are from auxiliary-field quantum Monte Carlo (AFQMC)Xu et al. (2024) and finite-temperature QMC (FTQMC) simulationsWhite et al. (1989) on the 2D Hubbard model. The inset of panel (a) compares the 4-orbital HK and QMC resultsXu et al. (2024) at $1/8$-hole-doping ($x=0.125$). LESW and DSWT in the exact solution of the 4-orbital HK model ($U=10$) for the hopping parameters shown in panels (b) and (c) respectively at $\beta=30$. The solid line shown with slope $2x$ is the band HK or atomic Hubbard result. The dashed and dotted lines depict the semiconductor and Fermi liquid results respectively. Note, there is no qualitative difference with the ED results for the Hubbard modelMeinders et al. (1993); Eskes et al. (1991). As $t/U$ increases, so does the DSWT.
@@
For any $n<N_{\rm sites}$, the orbital HK model corresponds to a $N_{\rm sites}/n$ copies of $n$-site Hubbard clusters all connected by TBC. The matrix that determines the TBC is purely local as it is determined by the dynamics. Hence, orbital HK is purely local in the rBZn regardless of the number of orbitals used, even $n=1$ corresponding to band HK whose locality is illustrated in Fig. 2. Theorem: The orbital HK model is equivalent to covering the Brillouin zone with $N_{\rm sites}/n$ Hubbard clusters each containing $n$-sites all connected via TBC. Clearly, the $N_{\rm sites}$-cluster Hubbard model results when $n=N_{\rm sites}$. Loosely, the $N_{\rm sites}/n$ identical copies throughout the Brillouin zone can be thought of as replica copies.
@@
Aside from DSWT, OHK also exhibits a pseudogap in the density of states. Shown in Fig. 6 is the density of states for the 4-orbital HK model with (Fig. 6(a,c)) and without (Fig. 6(b,d)) the next-nearest-neighbor hopping, $t^{\prime}$. Fig. 6(a) shows a suppression of the density of states at the chemical potential ($\rho(\omega=0)\lesssim 0.1$) in the underdoped region with $t^{\prime}=-0.25$ when $U\geq W$ ($W=8t$ is the bare bandwidth), indicative of a pseudogap in the absence of superconducting order. The complete tracking of the density of states at zero frequency in Fig. 6(c) displays the trend that as $U$ increases, the pseudo-gap region appears and extends to higher hole-doped density. In contrast, when $t^{\prime}=0$, the pseudo-gap suppression does not arise until $U\gg W$, as shown in Fig. 6(b,d). These observations are consistent with a recent studyHuang (2020) on a $2\times 2$ Hubbard cluster with TBC.
@@
Finally, we compute the heat capacity of the 4-orbital HK model at half-filling shown in Fig. 7. Most noticeable is the two-peak structure at $U>W$ representing a demarcation of the charge and spin degrees of freedom into high and low-temperature regimes, respectively, consistent with QMC simulations on the Hubbard model Duffy and Moreo (1997); Wang et al. (2022). Also at half-filling, we find a near-crossing of the heat capacity curves as a function of $U$ at a temperature intermediate between the spin and charge excitations. The Maxwell relations governing the entropy dictateVollhardt (1997) such a crossing. Since there is no sign-problem restriction, we can access the low-temperature heat capacity exactly. As we show in Fig. 7(b), the low-temperature heat capacity data follow a power-law increase detailed in the supplement. While such algebraic growth might seem counterintuitive for a Mott insulator, the first-excited state is charge neutralMai et al. (2024); Zhao et al. (2023b) as shown in the supplement. Such charge-neutral excitations determine the low-T behavior of the specific heat and are consistent with the absence of long-range magnetic order.
@@
IV Scaling Argument
The hidden power of the OHK model lies in its convergence properties in higher dimensions. This can be established entirely from the number of degrees of freedom that the orbital number introduces at each $k-$point. The quadratic dependence of the Hamiltonian on number operators implies that in 1d, there $n^{2}$ pairs of interactions at each k-point, hence the scaling $1/n^{2}$. In d-dimensions, the number of degrees is now $n^{2d}$ for $n^{d}$-orbital HK. Consequently, the convergence to Hubbard, guaranteed by Eq. (LABEL:ninfty), should scale as $1/n^{2d}$. This explains why so few orbitals are needed in 2d relative to 1d to match existing results. In the extreme limit of $d=\infty$, we find that this convergence factor vanishes for all $n>1$. Hence, $n=1$ or simple band HK is exact in $d=\infty$. This further explains why the spectral functions of band HK have a central peak as the Mott insulator is approached (see Fig. (2)) as in DMFTGeorges et al. (1996). Note focusing on the operator content is valid only if a fixed point exists. We have shown that HK represents a fixed pointZhao et al. (2023a); Huang et al. (2022b) as it contains the most relevant interaction with scaling dimension $-2$ thereby breaking the $Z_{2}$ symmetryZhao et al. (2023a); Huang et al. (2022b) at the Fermi surface of a Fermi liquid. Adding orbitals to each ${\bf k}$-point does not lead to further momentum integrations as such decorations (new degrees of freedom) of each ${\bf k}$-point just appear as discrete summations. Hence, the convergence to the fixed point should just be determined by the total number of degrees freedom at each k-point, namely $n^{2d}$. Consequently, the fixed pointZhao et al. (2023a); Huang et al. (2022b) controls Mott physics.
V Concluding Remarks:
The connection between OHK to Hubbard physics can also be understood in terms of momentum scattering. In the band HK model, the interaction is local in $k$ space and only zero momentum scattering/transfer obtains, while in the Hubbard model, the scattering momenta include all $q$ in the BZ. As we add orbitals to the band HK model, we effectively add additional momentum scattering. For example, the $2$-orbital HK includes $q=(\pi,\pi)$, and $4$-orbital HK includes $q=(\pi,\pi),(0,\pi),(\pi,0)$. A rigorous derivation to show this can be found in the supplement. That few-orbital HK already captures quantatively Hubbard physics supports the flow to Hubbard through a hierachy of momentum scattering.
The rapid convergence to Hubbard points to a new hybrid approach for OHK in higher dimensions: use DMRG for one of the directions and OHK on the others. Because of the rapid convergence of OHK, such an approach would essentially be exact. Consequently, OHK opens up a fundamentally new approach for simulations on strongly correlated electron matter as a result of the fixed pointZhao et al. (2023a). Moreover, given the consistent evolution of physical quantities with increasing orbital number, it is more amenable to machine learning technques than are other methods. This consistent evolution allows machine learning algorithms to detect patterns and make predictions more effectively, which could facilitate the discovery of new phases and transitions in these systems by leveraging the power of data-driven approaches.
@@ V Concluding Remarks:
. Acknowledgements We thank Barry Bradlyn for helpful comments and Shiwei Zhang for sending us his double occupancy data that is plotted in Fig. 5a (green dots) and George Sawatzky for a helpful e-mail exchange. This work was supported by the Center for Quantum Sensing and Quantum Materials, a DOE Energy Frontier Research Center, grant DE-SC0021238 (P. M. and P. W. P.). PWP also acknowledges NSF DMR-2111379 for partial funding of the HK work which led to these results.